OpenStudy (anonymous):

Integral:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \sin(x) }{ \sin(x)+\cos(x) }dx\]

OpenStudy (anonymous):

First I don't know where to begin I've tried substitution but I don't see it

OpenStudy (loser66):

oh, times both numerator and denominator by (sin - cos). Try it. I got the answer from there .

OpenStudy (anonymous):

Is it an Indefinite Integral or you have the bounds? something like:\[\int\limits_{0}^{\frac{\pi}{2}}\frac{ \sin x }{ \sin x+\cos x }dx\]

OpenStudy (anonymous):

@Loser66 you mean like multipling with conjugate

OpenStudy (anonymous):

@mukushla it's indefinite

OpenStudy (anonymous):

ok, go with @Loser66 then :-)

OpenStudy (loser66):

yes, and times (-1) to both to get cos(2x) from the denominator.

OpenStudy (anonymous):

@Loser66 continue

OpenStudy (anonymous):

@Loser66 continue please

OpenStudy (loser66):

hihihi... I did, on the paper.